# SpiderCat: A Catalog of Compact Binary Millisecond Pulsars

### v1

@@ 2.1 Sample
Figure 1: Histogram showing the discovery year of spider systems included in SpiderCat, color-coded by spider subclass. The discovery year is defined as the year in which a paper was published confirming the system as a spider or spider candidate, or reporting the detection of millisecond pulsations from the source. The first four spider systems are highlighted. The dramatic increase in discoveries following the launch of the Fermi Gamma-ray Space Telescope reflects its critical role in guiding targeted radio (e.g., 350-MHz GBT and 327-MHz Arecibo surveys) and multiwavelength surveys.
@@
Nr. Nr. $\gamma$-ray Nr. X-ray Nr. optical Nr. ATNF
RB 32 25 22 27 26
BW 49 34 21 14 39
RBc 23 21 10 21 0
@@
Total 109 84 55 67 82
At the time of writing, SpiderCat includes 109 systems: 32 confirmed RBs, 49 confirmed BWs, 23 RB candidates, and 5 BW candidates (see Table 1). We compiled all information from sources in the literature (as detailed in Table 2.2) and public catalogs, including the ATNF Pulsar Catalogue (R. N. Manchester et al., 2005) [^4] http://www.atnf.csiro.au/research/pulsar/psrcat/ and the Galactic MSP list (E. Ferrara & D. Lorimer, 2025) [^5] https://www.astro.umd.edu/~eferrara/GalacticMSPs.html, as well as several catalogs of multiwavelength data (see Section 3.5 for details).
Determining the precise discovery date of a spider system is often challenging, as many are found through radio surveys that require extensive post-observation analysis. To provide a consistent metric, we define the discovery year as the year in which a paper was published identifying the system as a spider or spider candidate, or reporting the detection of millisecond pulsations from the source. Fig. 1 shows a histogram of spider discoveries by year.
As seen in Fig. 1, the first known spider system, PSR B1957$+$20, was discovered in 1988 (A. S. Fruchter et al., 1988), with only three additional Galactic-field systems identified in the pre-Fermi era. Two of these were BWs, PSR J2051--0827 (B. W. Stappers et al., 1996) and PSR J0610--2100 (M. Burgay et al., 2006), while PSR J1723--2837 was discovered in 2004 (A. J. Faulkner et al., 2004) and later classified as a RB (F. Crawford et al., 2013). The three prominent spikes in the histogram correspond to two major radio surveys targeting unidentified Fermi sources; one with the GBT (six sources; J. W. T. Hessels et al., 2011), one with the Arecibo Observatory (five sources; H. T. Cromartie et al., 2016), and the more recent FAST-GPPS (13 sources; P. F. Wang et al., 2025).
2.2 Catalogue parameters
In Table 2.2, we list all sources along with the following parameters:
(1) Name – Source name. PSR if confirmed pulsar. Otherwise the name used in the discovery paper.
(2) Type – Spider type: BW = Black widow, RB = Redback, BWc = Black widow candidate, RBc = Redback candidate.
(3) Sub – Spider subtype: tr = Transitional millisecond pulsar, ti = Tidarren, hu = Huntsman.
(4) RA – Right ascension (J2000) (hh:mm:ss.s), where seconds are rounded to four digits.
(5) Dec – Declination (J2000) (±dd:mm:ss), where seconds are rounded to four digits.
(6) Year – Year of discovery publication.
(7) Ref. Disc. – Reference for the discovery publication.
(8) $P$ – Barycentric spin period of the pulsar (ms).
(9) $\dot{E}$ – Spin down luminosity (erg/s).
(10) $P_{\rm b}$ – Binary period of pulsar (hr).
(11) $M_{\rm c,min}$ – The minimum companion mass ($M_{\odot}$).
(12) Ref. Dyn. – Reference for the dynamical parameters (full timing solution when available).
@@ 2.2 Catalogue parameters
where $M_{\rm NS}$ is the mass of the pulsar, $M_{\rm c}$ is the mass of the companion, $i$ the orbital inclination, and $x_{\rm p}=a_{\rm p}\sin i/c$ is the projected semi-major axis of the pulsar’s orbit. Assuming $i=90^{\circ}$ and a “canonical” neutron star mass ($M_{\rm NS}=1.35$ M) yields a lower limit on the companion mass, denoted $M_{\rm c,min}$.
In addition, Section 3.5 presents the X-ray, $\gamma$-ray and optical counterparts when identified (see Tables 3 and 4), including the corresponding catalog source names and their angular separations from the spider coordinates listed in Table 2.2. Table 5 contains the derived source distances (described in detail in Section 3.4), as well as X-ray and $\gamma$-ray fluxes and luminosities, g-band apparent and absolute magnitudes, and g-r colors.
@@
Table 2: Spider catalogue. See Section 2.2 for details.
Name Type Sub RA DEC Year Ref P0 $\dot{E}$ Porb Mc,min Ref
(h:m:s) (d:m:s) Disc. (ms) ($10^{34}$ erg/s) (hr) (M) Dyn.
@@
PSR J0023+0923 BW 00:23:16.8803 +09:23:23.8902 2011 1 3.05 1.59 3.33 0.02 70
PSR J0251+2606 BW 02:51:2.5537 +26:06:9.97 2016 2 2.54 1.82 4.86 0.02 71
PSR J0312-0921 BW 03:12:6.2 -9:21:56 2021 3 3.70 1.53 2.34 0.01 16
PSR J0541+2959g BW 05:41:44 +29:59 2024 4 3.21 9.01 0.06 4
PSR J0610-2100 BW 06:10:13.5967 -21:00:27.8991 2006 5 3.86 0.85 6.86 0.02 72
PSR J0636+5128 BW ti 06:36:4.8471 +51:28:59.9658 2014 6 2.87 0.59 1.60 0.01 73
PSR J0952-0607 BW 09:52:8.3214 -06:07:23.49 2017 7 1.41 6.67 6.42 0.02 74
PSR J1124-3653 BW 11:24:1.116 -36:53:19.087 2011 1 2.41 1.70 5.45 0.03 75
PSR J1221-0633 BW 12:21:24.7582 -06:33:51.6998 2023 8 1.93 2.87 9.27 0.01 8
PSR J1301+0833 BW 13:01:38.26 +08:33:57.5 2012 9 1.84 6.65 6.54 0.02 16
PSR J1311-3430 BW ti 13:11:45.7242 -34:30:30.35 2012 10 2.56 4.93 1.56 0.01 76
PSR J1317-0157 BW 13:17:40.4489 -01:57:30.107 2023 8 2.91 0.88 2.14 0.02 8
PSR J1446-4701 BW 14:46:35.7092 -47:01:26.8138 2012 11 2.19 3.66 6.66 0.02 77
PSR J1513-2550 BW 15:13:23.3206 -25:50:31.285 2016 12 2.12 8.97 4.29 0.02 12
PSR J1544+4937 BW 15:44:4.4911 +49:37:55.374 2013 13 2.16 1.10 2.90 0.02 78
PSR J1555-2908 BW 15:55:40.6585 -29:08:28.4232 2022 14 1.79 30.78 5.60 0.05 14
PSR J1602-1009 BW 16:02:13.2805 -10:09:19.9199 2024 15 3.12 0.62 2.99 0.02 15
PSR J1627+3219 BW 16:27:52.9985 +32:18:26.643 2023 16 2.18 2.08 3.98 0.02 16
PSR J1630+3550 BW 16:30:35.949 +35:50:42.477 2022 17 3.23 2.64 7.58 0.01 53
PSR J1641+8049 BW 16:41:20.8331 +80:49:52.9233 2018 18 2.02 4.68 2.18 0.04 79
PSR J1653-0158 BW ti 16:53:38.0538 -01:58:36.893 2014 19 1.97 1.24 1.25 0.01 80
PSR J1705-1903 BW 17:05:43.8486 -19:03:41.4172 2019 20 2.48 4.41 0.04 77
PSR J1720-0533 BW 17:20:54.5059 -05:34:23.8223 2021 21 3.27 0.92 3.16 0.03 81
PSR J1731-1847 BW 17:31:17.6098 -18:47:32.666 2011 22 2.34 7.78 7.47 0.03 82
PSR J1745-23 BW 17:45:30 -23:25 2020 23 5.42 3.97 0.03 23
PSR J1745+1017 BW 17:45:33.8371 +10:17:52.523 2013 24 2.65 0.58 17.53 0.01 24
PSR J1805+0615 BW 18:05:42.3997 +06:15:18.606 2016 2 2.13 9.31 8.08 0.02 71
PSR J1810+1744 BW 18:10:37.2848 +17:44:37.38 2011 1 1.66 3.84 3.56 0.04 75
PSR J1814+0045g BW 18:14:10 +00:45 2024 4 2.31 4.86 0.06 4
PSR J1830-0106g BW 18:30:7 -01:06 2024 4 1.76 2.56 0.05 4
PSR J1833-3840 BW 18:33:4.6 -38:40:46 2023 16 1.87 10.77 21.61 0.01 83
PSR J1838+1507g BW 18:38:36 +15:07 2024 4 3.82 2.69 0.03 4
PSR J1847+0342g BW 18:47:30 +03:42 2024 4 4.29 3.34 0.02 4
PSR J1919+0126g BW 19:19:23 +01:26 2024 4 1.90 6.47 0.02 4
PSR J1928+1245 BW 19:28:45.3936 +12:45:53.374 2019 25 3.02 2.40 3.28 0.01 25
PSR J1946-5403 BW 19:46:34.497 -54:03:42.51 2015 26 2.71 3.12 0.02 26
PSR J1953+1006g BW 19:53:34 +10:06 2024 4 2.59 2.89 0.01 4
PSR B1957+20 BW 19:59:36.7699 +20:48:15.1222 1988 27 1.61 16.02 9.17 0.02 84
@@
References. — 1: J. W. T. Hessels et al. (2011) 2: H. T. Cromartie et al. (2016) 3: S. Tabassum et al. (2021) 4: P. F. Wang et al. (2025) 5: M. Burgay et al. (2006) 6: K. Stovall et al. (2014) 7: C. G. Bassa et al. (2017) 8: J. K. Swiggum et al. (2023) 9: P. S. Ray et al. (2012) 10: R. W. Romani (2012) 11: M. J. Keith et al. (2012) 12: S. Sanpa-arsa (2016) 13: B. Bhattacharyya et al. (2013) 14: P. S. Ray et al. (2022) 15: A. E. McEwen et al. (2024) 16: D. A. Smith et al. (2023) 17: C. Sobey et al. (2022) 18: R. S. Lynch et al. (2018) 19: R. W. Romani et al. (2014) 20: V. Morello et al. (2019) 21: S. Q. Wang et al. (2021) 22: S. D. Bates et al. (2011) 23: A. D. Cameron et al. (2020) 24: E. D. Barr et al. (2013b) 25: E. Parent et al. (2019) 26: F. Camilo et al. (2015) 27: A. S. Fruchter et al. (1988) 28: B. W. Stappers et al. (1996) 29: L. Guillemot et al. (2019) 30: S. M. Ransom et al. (2011) 31: M. J. Keith et al. (2011) 32: J. Boyles et al. (2011) 33: K.-L. Li et al. (2021) 34: K. B. Burdge et al. (2022) 35: S. J. Swihart et al. (2022) 36: A. V. Karpova et al. (2025) 37: D. A. Zyuzin et al. (2024) 38: K.-L. Li et al. (2016) 39: J. P. Halpern et al. (2017) 40: K.-L. Li et al. (2018) 41: A. M. Archibald et al. (2009) 42: C. J. Clark et al. (2023b) 43: C. G. Bassa et al. (2014) 44: A. Ghosh et al. (2024) 45: E. F. Keane et al. (2018) 46: F. Camilo et al. (2016) 47: S. D. Bates et al. (2015) 48: A. J. Faulkner et al. (2004) 49: K.-Y. Au et al. (2023) 50: J. Strader et al. (2025) 51: P. Lazarus et al. (2015) 52: R. W. Romani (2015) 53: E. F. Lewis et al. (2023) 54: S. J. Swihart et al. (2020) 55: R. W. Romani & M. S. Shaw (2011b) 56: C. Lu et al. (2025) 57: J. M. Miller et al. (2020) 58: J. Strader et al. (2016) 59: J. Strader et al. (2014) 60: R. Kyer et al. (2025) 61: M. Turchetta et al. (2024) 62: S. J. Swihart et al. (2017) 63: Z. Wang et al. (2020) 64: S. J. Swihart et al. (2021) 65: F. Coti Zelati et al. (2019) 66: S. Bogdanov & J. P. Halpern (2015) 67: A. Zic et al. (2024) 68: R. H. D. Corbet et al. (2022) 69: A. V. Karpova et al. (2023) 70: M. F. Alam et al. (2021a) 71: J. S. Deneva et al. (2021) 72: E. van der Wateren et al. (2022) 73: W. Fiore et al. (2023) 74: L. Nieder et al. (2019) 75: P. Bangale et al. (2024) 76: H. J. Pletsch et al. (2012) 77: M. Shamohammadi et al. (2024) 78: S. Kumari et al. (2023) 79: A. Y. Kirichenko et al. (2024) 80: L. Nieder et al. (2020) 81: C.-C. Miao et al. (2023) 82: C. Ng et al. (2014) 83: M. Kerr et al. (2025) 84: Z. Arzoumanian et al. (1994) 85: G. Shaifullah et al. (2016) 86: M. F. Alam et al. (2021b) 87: D. J. Reardon et al. (2021) 88: K. Crowter et al. (2020) 89: K. I. Perez et al. (2023) 90: T. Thongmeearkom et al. (2024) 91: A. M. Archibald et al. (2013) 92: M. Burgay et al. (2024) 93: J. Roy et al. (2015) 94: F. Crawford et al. (2013) 95: O. G. Dodge et al. (2024) 96: K. Stovall et al. (2016) 97: C. J. Clark et al. (2021) 98: H. J. Pletsch & C. J. Clark (2015)

### v4

@@ 2.1 Sample
No. No. $\gamma$ ray No. X-Ray No. Optical No. ATNF
RB 30 23 21 25 25
BW 50 35 20 19 39
HM 2 2 1 2 1
RBc 23 22 10 21 0
@@
HMc 1 0 1 1 0
Total 111 86 55 73 65
At the time of writing, SpiderCat includes 111 systems: 30 confirmed RBs, 50 confirmed BWs, two confirmed HMs, 23 RBcs, five BWcs, and one HM candidate (HMc; see Table 1). We compiled all information from sources in the literature (as detailed in Table 2) and public catalogs, including the ATNF Pulsar Catalog (R. N. Manchester et al., 2005) [^5] http://www.atnf.csiro.au/research/pulsar/psrcat/ and the Galactic MSP list (E. Ferrara & D. Lorimer, 2025), [^6] https://www.astro.umd.edu/~eferrara/GalacticMSPs.html as well as several catalogs of multiwavelength data (see Section 3.5 for details).
Determining the precise discovery date of a spider system is often challenging, as many are found through radio surveys that require extensive postobservation analysis. To provide a consistent metric, we define the discovery year as the year in which a paper was published identifying the system as a spider or spider candidate, or reporting the detection of millisecond pulsations from the source. Figure 1 shows a histogram of spider discoveries by year.
As seen in Figure 1, the first known spider system, PSR B1957++20, was discovered in 1988 (A. S. Fruchter et al., 1988), with only three additional Galactic-field systems identified in the pre-Fermi era. Two of these were BWs, PSR J2051-0827 (B. W. Stappers et al., 1996) and PSR J0610-2100 (M. Burgay et al., 2006), while PSR J1723-2837 was discovered in 2004 (A. J. Faulkner et al., 2004) and later classified as an RB (F. Crawford et al., 2013). The three prominent spikes in the histogram correspond to two major radio surveys targeting unidentified Fermi sources: one with the GBT (six sources; J. W. T. Hessels et al., 2011), one with the Arecibo Observatory (five sources; H. T. Cromartie et al., 2016), and the more recent FAST-GPPS (13 sources; P. F. Wang et al., 2025).
2.2 Catalog Parameters
In Table 2, we list all sources along with the following parameters:
(1) Name—source name. “PSR” if confirmed pulsar. Otherwise the name used in the discovery paper.
(2) Type—spider type: BW, RB, HM, BWc, RBc, and HMc.
(3) Sub.—spider subtype: tr = tMSP, ti = tidarren.
(4) R.A.—(J2000) (hh:mm:ss.ssss), where seconds are rounded to four digits.
(5) Decl.—(J2000) (±dd:mm:ssss), where seconds are rounded to four digits.
(6) Year—year of discovery publication.
(7) Ref. discovery—reference for the discovery publication.
(8) PP—barycentric spin period of the pulsar (milliseconds).
(9) $\dot{E}$—spin-down luminosity (in units of erg s-1).
(10) $P_{\rm b}$—binary period of pulsar (hours).
(11) $M_{\rm c,min}$—the minimum companion mass ($M_{\odot}$).
(12) Ref. dyn.—reference for the dynamical parameters (full timing solution when available).
@@ 2.2 Catalog Parameters
where $M_{\rm NS}$ is the mass of the pulsar, $M_{\rm c}$ is the mass of the companion, $i$ the orbital inclination, and $x_{\rm p}=a_{\rm p}\sin i/c$ is the projected semimajor axis of the pulsar’s orbit. Assuming $i=90^{\circ}$ and a “canonical” neutron star mass ($M_{\rm NS}=1.35$ M) yields a lower limit on the companion mass, denoted $M_{\rm c,min}$.
In addition, Section 3.5 presents the X-ray, $\gamma$-ray and optical counterparts when identified (see Tables 3 and 4), including the corresponding catalog source names and their angular separations from the spider coordinates listed in Table 2. Table 5 contains the derived source distances (described in detail in Section 3.4), as well as X-ray and $\gamma$-ray fluxes and luminosities, $g$-band apparent and absolute magnitudes, and grg-r colors.
@@
Table 2: Spider Catalog footnotetext: Notes. See Section 2.2 for details. (This table is available in machine-readable form in the online article.)
Name Type Sub. R.A. Decl. Year Ref. P $\dot{E}$ Pb Mc,min Ref.
($h$:$m$:$s$.ssss) ($d$:$m$:$s$.ssss) Discovery (ms) ($10^{34}$ erg s-1) (hr) (M) Dyn.
@@
PSR J0023+0923 BW 00:23:16.8803 +09:23:23.8902 2011 1 3.05 1.59 3.33 0.016 71
PSR J0251+2606 BW 02:51:2.5537 +26:06:9.97 2016 2 2.54 1.82 4.86 0.024 72
PSR J0312-0921 BW 03:12:6.2147 -09:21:56.5532 2021 3 3.70 1.53 2.34 0.009 16
PSR J0541+2959g BW 05:41:44 +29:59 2024 4 3.21 9.01 0.057 4
PSR J0610-2100 BW 06:10:13.5967 -21:00:27.8991 2006 5 3.86 0.85 6.86 0.021 73
PSR J0636+5128 BW ti 06:36:4.8471 +51:28:59.9658 2014 6 2.87 0.59 1.60 0.007 74
PSR J0952-0607 BW 09:52:8.3214 -06:07:23.49 2017 7 1.41 6.67 6.42 0.019 75
PSR J1124-3653 BW 11:24:1.116 -36:53:19.087 2011 1 2.41 1.7 5.45 0.027 76
PSR J1221-0633 BW 12:21:24.7582 -06:33:51.6998 2023 8 1.93 2.87 9.27 0.013 8
PSR J1301+0833 BW 13:01:38.26 +08:33:57.5 2012 9 1.84 6.65 6.54 0.024 16
PSR J1311-3430 BW ti 13:11:45.7242 -34:30:30.35 2012 10 2.56 4.93 1.56 0.008 77
PSR J1317-0157 BW 13:17:40.4489 -01:57:30.107 2023 8 2.91 0.88 2.14 0.018 8
PSR J1356+0230aaThis pulsar was discovered by TRAPUM’s Fermi Sources working group using MeerKAT, targeting the sky position of 4FGL J1356.6+0234. The pulsar has been timed in $\gamma$-rays by Paolo Freire; https://confluence.slac.stanford.edu/spaces/GLAMCOG/pages/383919352/Gamma-ray+PSR+J1356p0230. BW 13:56:37.2 +02:30:28.8 2021 2.83 1.37
PSR J1446-4701 BW 14:46:35.7092 -47:01:26.8138 2012 11 2.19 3.66 6.66 0.019 78
PSR J1513-2550 BW 15:13:23.3206 -25:50:31.285 2016 12 2.12 8.97 4.29 0.016 12
PSR J1544+4937 BW 15:44:4.4911 +49:37:55.374 2013 13 2.16 1.1 2.90 0.017 79
PSR J1555-2908 BW 15:55:40.6585 -29:08:28.4232 2022 14 1.79 30.8 5.60 0.051 14
PSR J1602-1009 BW 16:02:13.2805 -10:09:19.9199 2024 15 3.12 0.62 2.99 0.018 15
PSR J1627+3219 BW 16:27:52.9985 +32:18:26.643 2023 16 2.18 2.08 3.98 0.022 16
PSR J1630+3550 BW 16:30:35.949 +35:50:42.477 2022 17 3.23 2.64 7.58 0.01 54
PSR J1641+8049 BW 16:41:20.8331 +80:49:52.9233 2018 18 2.02 4.68 2.18 0.04 80
PSR J1653-0158 BW ti 16:53:38.0538 -01:58:36.893 2014 19 1.97 1.24 1.25 0.01 81
PSR J1705-1903 BW 17:05:43.8486 -19:03:41.4172 2019 20 2.48 4.41 0.041 78
PSR J1720-0533 BW 17:20:54.5059 -05:34:23.8223 2021 21 3.27 0.93 3.16 0.029 82
PSR J1731-1847 BW 17:31:17.6098 -18:47:32.666 2011 22 2.34 7.78 7.47 0.033 83
PSR J1745-23 BW 17:45:30 -23:25 2020 23 5.42 3.97 0.026 23
PSR J1745+1017 BW 17:45:33.8371 +10:17:52.523 2013 24 2.65 0.58 17.53 0.014 24
PSR J1805+0615 BW 18:05:42.3997 +06:15:18.606 2016 2 2.13 9.31 8.08 0.023 72
PSR J1810+1744 BW 18:10:37.2848 +17:44:37.38 2011 1 1.66 3.84 3.56 0.043 76
PSR J1814+0045g BW 18:14:10 +00:45 2024 4 2.31 4.86 0.055 4
PSR J1830-0106g BW 18:30:7 -01:06 2024 4 1.76 2.56 0.046 4
PSR J1833-3840 BW 18:33:4.5823 -38:40:46.072 2023 16 1.87 10.8 21.61 0.008 84
PSR J1838+1507g BW 18:38:36 +15:07 2024 4 3.82 2.69 0.028 4
PSR J1847+0342g BW 18:47:30 +03:42 2024 4 4.29 3.34 0.017 4
PSR J1919+0126g BW 19:19:23 +01:26 2024 4 1.90 6.47 0.025 4
PSR J1928+1245 BW 19:28:45.3936 +12:45:53.374 2019 25 3.02 2.4 3.28 0.009 25
PSR J1946-5403 BW 19:46:34.497 -54:03:42.51 2015 26 2.71 3.12 0.021 26
PSR J1953+1006g BW 19:53:34 +10:06 2024 4 2.59 2.89 0.012 4
PSR B1957+20 BW 19:59:36.7699 +20:48:15.1222 1988 27 1.61 16.0 9.17 0.021 85
@@
References. — (1) J. W. T. Hessels et al. (2011); (2) H. T. Cromartie et al. (2016); (3) S. Tabassum et al. (2021); (4) P. F. Wang et al. (2025); (5) M. Burgay et al. (2006); (6) K. Stovall et al. (2014); (7) C. G. Bassa et al. (2017); (8) J. K. Swiggum et al. (2023); (9) P. S. Ray et al. (2012); (10) R. W. Romani (2012); (11) M. J. Keith et al. (2012); (12) S. Sanpa-arsa (2016); (13) B. Bhattacharyya et al. (2013); (14) P. S. Ray et al. (2022); (15) A. E. McEwen et al. (2024); (16) D. A. Smith et al. (2023); (17) C. Sobey et al. (2022); (18) R. S. Lynch et al. (2018); (19) R. W. Romani et al. (2014); (20) V. Morello et al. (2019); (21) S. Q. Wang et al. (2021); (22) S. D. Bates et al. (2011); (23) A. D. Cameron et al. (2020); (24) E. D. Barr et al. (2013); (25) E. Parent et al. (2019); (26) F. Camilo et al. (2015); (27) A. S. Fruchter et al. (1988); (28) B. W. Stappers et al. (1996); (29) L. Guillemot et al. (2019); (30) S. M. Ransom et al. (2011); (31) M. J. Keith et al. (2011); (32) J. Boyles et al. (2011); (33) K.-L. Li et al. (2021); (34) K. B. Burdge et al. (2022); (35) S. J. Swihart et al. (2022); (36) A. V. Karpova et al. (2025); (37) D. A. Zyuzin et al. (2024); (38) F. Camilo et al. (2016); (39) J. Strader et al. (2025); (40) S. J. Swihart et al. (2017); (41) K.-L. Li et al. (2016); (42) J. P. Halpern et al. (2017); (43) K.-L. Li et al. (2018); (44) A. M. Archibald et al. (2009); (45) C. J. Clark et al. (2023b); (46) C. G. Bassa et al. (2014); (47) A. Ghosh et al. (2024); (48) E. F. Keane et al. (2018); (49) S. D. Bates et al. (2015); (50) A. J. Faulkner et al. (2004); (51) K.-Y. Au et al. (2023); (52) P. Lazarus et al. (2015); (53) R. W. Romani (2015); (54) E. F. Lewis et al. (2023); (55) S. J. Swihart et al. (2020); (56) R. W. Romani & M. S. Shaw (2011); (57) C. Lu et al. (2025); (58) J. M. Miller et al. (2020); (59) J. Strader et al. (2016); (60) J. Strader et al. (2014); (61) J. Strader et al. (2021); (62) R. Kyer et al. (2025); (63) M. Turchetta et al. (2024); (64) Z. Wang et al. (2020); (65) S. J. Swihart et al. (2021); (66) F. Coti Zelati et al. (2019); (67) S. Bogdanov & J. P. Halpern (2015); (68) A. Zic et al. (2024); (69) R. H. D. Corbet et al. (2022); (70) A. V. Karpova et al. (2023); (71) M. F. Alam et al. (2021a); (72) J. S. Deneva et al. (2021); (73) E. van der Wateren et al. (2022); (74) W. Fiore et al. (2023); (75) L. Nieder et al. (2019); (76) P. Bangale et al. (2024); (77) H. J. Pletsch et al. (2012); (78) M. Shamohammadi et al. (2024); (79) S. Kumari et al. (2023); (80) A. Y. Kirichenko et al. (2024); (81) L. Nieder et al. (2020); (82) C.-C. Miao et al. (2023); (83) C. Ng et al. (2014); (84) M. Kerr et al. (2025); (85) Z. Arzoumanian et al. (1994); (86) G. Shaifullah et al. (2016); (87) M. F. Alam et al. (2021b); (88) D. J. Reardon et al. (2021); (89) K. Crowter et al. (2020); (90) K. I. Perez et al. (2023); (91) T. Thongmeearkom et al. (2024); (92) A. M. Archibald et al. (2013); (93) M. Burgay et al. (2024); (94) J. Roy et al. (2015); (95) F. Crawford et al. (2013); (96) O. G. Dodge et al. (2024); (97) K. Stovall et al. (2016); (98) C. J. Clark et al. (2021); (99) H. J. Pletsch & C. J. Clark (2015)